A man travels from A to B at 4 km/h and from B to A at 6 km/h. The total journey takes 45 minutes. Find the distance travelled.
Answer by mananth(16946)

let the distance between A & Bbe x miles
time = d/r
Time A to B + time B to A = 3/4 hours
x/4 +x/6 = 3/4
LCD = 12
multiply equation by 12
3x+2x=9
5x=9
x= 9/5 miles